OpenStudy (anonymous):

Could someone give me a quick lesson about absolute value equations & use this problem? reward will be given. Solve for x: −5|x + 1| = 10

OpenStudy (anonymous):

|-5x+5|=10 ( -5x+5)^{2}=10^{2}/ 25x^2-50x+25=100 25x^2-50x-75=0 x^2-2x-3=0 (x-3)(x+1) x=3 or x=-1